### v1

# Universal Reverse Information Projections and Optimal E-statistics

Abstract
Information projections have found many important applications in probability theory, statistics, and related fields. In the field of hypothesis testing in particular, the reverse information projection (RIPr) has recently been shown to lead to so-called growth-rate optimal (GRO) e-statistics for testing simple alternatives against composite null hypotheses. However, the RIPr as well as the GRO criterion are only defined in cases where the infimum information divergence between the null and alternative is finite. Here, we show that under much weaker conditions there often still exists an element in the alternative that is ‘closest’ to the null: the universal reverse information projection. The universal reverse information projection and its non-universal counterpart coincide whenever the KL is finite, and the strictness of this generalization will be shown by an example. Furthermore, the universal RIPr leads to optimal e-statistics in a sense that is a novel, but natural, extension of the GRO criterion. Finally, we discuss conditions under which the universal RIPr is a strict sub-probability distributions, and conditions under which an approximation of the universal RIPr leads to approximate e-statistics.
I Introduction
We write $D(\nu\|\lambda)$ for the information divergence (Kullback-Leibler divergence, [1, 2, 3]) between two finite measures $\nu$ and $\lambda$ given by
$D(\nu\|\lambda)=\begin{cases}\displaystyle\int_{\Omega}\ln\left(\frac{\mathrm{d}\nu}{\mathrm{d}\lambda}\right)\,\mathrm{d}\nu-(\nu(\Omega)-\lambda(\Omega)),&\text{if }\nu\ll\lambda;\\ \infty,&\text{else.}\end{cases}$
For probability measures the interpretation of $D(\nu\|\lambda)$ is that it measures how much we gain by coding according to $\nu$ rather than coding according to $\lambda$ if data are distributed according to $\nu$. Many problems in probability theory and statistics, such as conditioning and maximum likelihood estimation, can be cast as minimization in either or both arguments of the information divergence. In particular, this is the case within the recently established and now flourishing theory of hypothesis testing based on e-statistics that allows for optional continuation of experiments (see Section II-C)[4, 5]. That is, a kind of duality has been established between optimal e-statistics for testing a simple alternative $P$ against a composite null hypothesis $\mathcal{C}$ and reverse information projections [4]. Here, the reverse information projection (RIPr) of $P$ on $\mathcal{C}$ is — if it exists — a unique measure $\hat{Q}$ such that every sequence $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ with $D(P\|Q_{n})\to\inf_{Q\in\mathcal{C}}D(P\|Q)$ converges to $\hat{Q}$ in a particular way [6, 7]. It has been shown that whenever $\mathcal{C}$ is convex and $D(P\|\mathcal{C}):=\inf_{Q\in\mathcal{C}}D(P\|Q)<\infty$, the RIPr $\hat{Q}$ exists and the likelihood ratio between $P$ and $\hat{Q}$ is the optimal e-statistic for testing $P$ against $\mathcal{C}$. However, it is clear that the RIPr does not exist if the information divergence between $P$ and $\mathcal{C}$ is infinite, i.e. $D(P\|\mathcal{C})=\infty$. This leaves a void in the theory of optimality of e-statistics. In this work we remedy this by realizing that even if all measures in $\mathcal{C}$ are infinitely worse than $P$ at describing data distributed according to $P$ itself, there can still be a measure that performs best relative to the elements of $\mathcal{C}$. To find such a measure, we consider the description gain [8] given by
$D(P\|Q\rightsquigarrow Q^{\prime})=\int_{\Omega}\ln\left(\frac{\mathrm{d}Q^{\prime}}{\mathrm{d}Q}\right)\,\mathrm{d}P-(Q^{\prime}(\Omega)-Q(\Omega))$ (1)
whenever this integral is well-defined. If the quantities involved are finite then the description gain reduces to
$D(P\|Q\rightsquigarrow Q^{\prime})=D(P\|Q)-D(P\|Q^{\prime}).$ (2)
In analogy to the interpretation of information divergence for coding, description gain measures how much we gain by coding according to $Q^{\prime}$ rather than $Q$ if data are distributed according to $P$. Furthermore denote
$D(P\|Q\rightsquigarrow\mathcal{C}):=\sup_{Q^{\prime}\in\mathcal{C}}D(P\|Q\rightsquigarrow Q^{\prime}),$
where undefined values are counted as $\infty$- ∞ when taking the supremum. If there exists at least one $Q^{*}\in\mathcal{C}$ such that $P\ll Q^{*}$, then $D(P\|Q\rightsquigarrow\mathcal{C})$ is a well-defined number in $[0,\infty]$ for any $Q\in\mathcal{C}$. This quantity should be seen as the maximum description gain one can get by switching from $Q$ to any other measure in $\mathcal{C}$.
Intuitively, if there is a best descriptor in $\mathcal{C}$, nothing can be gained by switching away from it. Indeed, in Proposition 2 we show that $\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})$ is finite if and only if it is equal to zero. Furthermore, in Theorem 3 we show that — under very mild conditions — there exists a unique measure $\hat{Q}$ such that every sequence $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ with
$D(P\|Q_{n}\rightsquigarrow\mathcal{C})\to 0$
converges to $\hat{Q}$ in a specific way. We call $\hat{Q}$ the universal RIPr, as it coincides with the RIPr whenever the information divergence is finite. Furthermore, in Theorem 5 we show that whenever the universal RIPr $\hat{Q}$ exists, the likelihood ratio of $P$ and $\hat{Q}$ is an optimal e-statistic in a sense that can be seen as a strict generalization of previously known optimality criteria for e-statistics. Finally, Theorem 4 and proposition 4 provide certain properties of the universal RIPr that give insights even in the finite information divergence setting.
II Background
II-A Preliminaries
We work with a measurable space $(\Omega,\mathcal{F})$ and, unless specified otherwise, all measures will be defined on this space. Throughout, $P$ will denote a finite measure and $\mathcal{C}$ a set of finite measures, such that $P$ and all $Q\in\mathcal{C}$ have densities w.r.t. a common measure $\mu$. These densities will be denoted with lowercase, i.e. $p$ and $q$ respectively. We will assume throughout that $\mathcal{C}$ is $\sigma$-convex, i.e. closed under countable mixtures, though we will refer to this simply as ‘convex’. Furthermore, we assume that there exists at least one $Q^{*}\in\mathcal{C}$ such that $P\ll Q^{*}$. On the one hand, this ensures that $D(P\|Q\rightsquigarrow\mathcal{C})$ is a well-defined number in $[0,\infty]$ for any $Q\in\mathcal{C}$. On the other hand, it aligns with our philosophy when we turn to hypothesis testing, in which case $P$ and all $Q\in\mathcal{C}$ will be probability distributions and serve as the alternative and null hypothesis respectively. We will consider $P$ mostly as a tool to gather evidence against $\mathcal{C}$, so that it does not make sense to consider the case in which $P$ puts mass on events that cannot occur according to the null, as the null hypothesis can be discredited in such scenarios regardless of how much mass $P$ puts on the event.
II-B The Reverse Information Projection
As mentioned briefly above, the reverse information projection is the result of minimizing the information divergence between $P$ and $\mathcal{C}$. If $\mathcal{C}$ is an exponential family, this problem is well understood [9], but we focus here on the case that $\mathcal{C}$ is a general convex set. In this setting, the following theorem establishes existence and uniqueness of a limiting object for any sequence $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ such that $D(P\|Q_{n})\to D(P\|\mathcal{C})$ whenever the latter is finite. This limit (i.e. $\hat{Q}$ in the following) is called the reverse information projection of $P$ on $\mathcal{C}$.
Theorem 1 (Li [6], Definition 4.2 and Theorem 4.3).
If $P$ and all $Q\in\mathcal{C}$ are probability distributions such that $D(P\|\mathcal{C})<\infty$, then there exists a unique (potentially sub-) probability distribution $\hat{Q}$ such that:
  1. 1.
    We have that $\ln q_{n}\rightarrow\ln\hat{q}$ in $L_{1}(P)$ for all sequences $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ such that $\lim_{n\rightarrow\infty}D(P\|Q_{n})=D(P\|\mathcal{C})$.
  2. 2.
    $\int_{\Omega}\ln\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}=D(P\|\mathcal{C})$,
  3. 3.
    $\int_{\Omega}\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}Q\leq 1$ for all $Q\in\mathcal{C}$.
II-C E-statistics and Growth Rate Optimality
The e-value has recently emerged as a popular alternative to the p-value for hypothesis testing [5, 10, 11]. It can be thought of as a measure of statistical evidence that is intimately linked with numerous ideas, such as likelihood ratios, test martingales [12] and tests of randomness [13]. Formally, an e-value is defined as the value taken by an e-statistic, which is defined as a random variable $E:\Omega\rightarrow[0,\infty]$ that satisfies $\int_{\Omega}E\,\mathrm{d}Q\leq 1$ for all $Q\in\mathcal{C}$ [14]. The set of all e-statistics is denoted as $\mathcal{E}_{\mathcal{C}}$. Large e-values constitute evidence against $\mathcal{C}$ as null hypothesis, so that the null can be rejected when the computed e-value exceeds a certain threshold. For example, the test that rejects the null hypothesis when $E\geq\nicefrac{{1}}{{\alpha}}$ has a type-I error guarantee of $\alpha$ by a simple application of Markov’s inequality: $Q(E\geq\nicefrac{{1}}{{\alpha}})\leq\alpha\int_{\Omega}E\,\mathrm{d}Q\leq\alpha$.
In general, the set $\mathcal{E}_{\mathcal{C}}$ of e-statistics is quite large, and the above does not tell us which e-statistic to pick. This question was studied in [4] and a log-optimality criterion coined GRO was introduced for the case that the interest is in gaining as much evidence as possible relative to an alternative hypothesis given by a single probability distribution $P$. This criterion can be traced back to the Kelly betting criterion in [15] and is further discussed in [16].
Definition 1.
If it exists, an e-statistic $\hat{E}\in\mathcal{E}_{\mathcal{C}}$ is Growth-Rate Optimal (GRO) if it achieves
$\int_{\Omega}\ln\hat{E}\,\mathrm{d}P=\sup_{E\in\mathcal{E}_{\mathcal{C}}}\int_{\Omega}\ln E\,\mathrm{d}P.$
The following theorem establishes a duality between GRO e-statistics and reverse information projections. For a limited set of testing problems, it states that GRO e-statistics exist and are uniquely given by likelihood ratios.
Theorem 2 (Grünwald et al. [4], Theorem 1).
If $P$ and all $Q\in\mathcal{C}$ are probability distributions such that $D(P\|\mathcal{C})<\infty$, $p(\omega)>0$ for all $\omega\in\Omega$, and $\hat{Q}$ is the RIPr of $P$ on $\mathcal{C}$, then $\hat{E}=\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}$ is GRO with rate equal to $D(P\|\mathcal{C})$, i.e.
$\sup_{E\in\mathcal{E}_{\mathcal{C}}}\int_{\Omega}\ln E\,\mathrm{d}P=\int_{\Omega}\ln\hat{E}\,\mathrm{d}P=D(P\|\mathcal{C}).$
Furthermore, for any GRO e-statistic $\tilde{E}$, we have that $\tilde{E}=\hat{E}$ holds $P$-almost surely.
III The Universal Reverse Information Projection
In this section, we will prove a result analogous to Theorem 1 in a more general setting. Rather than convergence of the logarithm of densities in $L_{1}(P)$, we consider convergence with respect to a metric on the set of measurable positive functions, i.e. $\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)=\{f:\Omega\to\mathbb{R}_{>0}:f\text{ measurable}\}$. For $f,f^{\prime}\in\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$ we define
$m_{P}^{2}(f,f^{\prime}):=\frac{1}{2}\int_{\Omega}\ln\left(\frac{\overline{f}}{f}\right)+\ln\left(\frac{\overline{f}}{f^{\prime}}\right)\,\mathrm{d}P,$ (3)
where $\overline{f}:=\nicefrac{{(f+f^{\prime})}}{{2}}$. This is a divergence that can be thought of as the averaged Bregman divergence associated with the convex function $\gamma(x)=x-1-\ln(x)$. In [17], such divergences are studied in detail for general $\gamma$. In particular, they show that the function
$m^{2}_{\gamma}(x,y)=\frac{1}{2}\gamma(x)+\frac{1}{2}\gamma(y)-\gamma\left(\frac{x+y}{2}\right)$
is the square of a metric if and only if $\ln\left(\gamma^{\prime\prime}(x)\right)^{\prime\prime}\geq 0$. In our case, $\ln(\gamma^{\prime\prime}(x))^{\prime\prime}=2x^{-2}$, so this result holds. This can be used together with an application of Minkowski inequality to show that the triangle inequality holds for the square root of the divergence (3), i.e. $m_{P}$, on $\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$. It should also be clear that for $f,g\in\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$ if $f=g$ everywhere, then $m_{P}(f,g)=0$. Conversely $m_{P}(f,g)=0$ only implies that $P(f\neq g)=0$. This prevents us from calling $m_{P}$ a metric on $\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$, and we therefore define, analogous to $\mathcal{L}^{p}$ and $L^{p}$ spaces, $M\left(\Omega,\mathbb{R}_{>0}\right)$ as the set of equivalence classes of $\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$ under the relation $\sim$’ given by $f\sim g\Leftrightarrow P(f\neq g)=0$. By the discussion above, $m_{P}$ properly defines a metric on $M\left(\Omega,\mathbb{R}_{>0}\right)$. In the following we will often ignore this technicality and simply act as if $m_{P}$ defines a metric on $\mathcal{M}\left(\Omega,\mathbb{R}_{>0}\right)$, since we are not interested in what happens on null sets of $P$.
Proposition 1.
The metric space $(M\left(\Omega,\mathbb{R}_{>0}\right),m_{P})$ is complete.
Everything is now in place to state the main result.
Theorem 3.
If $\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})<\infty$, then there exists a measure $\hat{Q}$ that satisfies the following for every sequence $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ s.t. $D(P\|Q_{n}\rightsquigarrow\mathcal{C})\to\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})$ for $n\to\infty$:
  1. 1.
    $q_{n}\rightarrow\hat{q}$ in $m_{P}$,
  2. 2.
    If $P^{\prime}$ is a measure such that $\left|\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow P^{\prime})\right|<\infty$ then $\int_{\Omega}\ln\frac{\mathrm{d}P^{\prime}}{\mathrm{d}\hat{Q}}\,\mathrm{d}P=\lim_{n\to\infty}\int_{\Omega}\ln\frac{\mathrm{d}P^{\prime}}{\mathrm{d}Q_{n}}\,\mathrm{d}P$,
  3. 3.
    $\int_{\Omega}\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}Q\leq P(\Omega)+Q(\Omega)-\liminf_{n\to\infty}Q_{n}(\Omega)$ for any $Q\in\mathcal{C}$.
Theorem 1 is a special case of Theorem 3 when $P$ and all $Q\in\mathcal{C}$ are probability distributions and $D(P\|\mathcal{C})<\infty$. This follows because Equation (2) implies that minimizing $D(P\|Q\rightsquigarrow Q^{\prime})$ over $Q$ is equivalent to minimizing $D(P\|Q)$. The measure $\hat{Q}$ as in Theorem 3 therefore extends the notion of the reverse information projection of $P$ on $\mathcal{C}$. We call $\hat{Q}$ the universal reverse information projection of $P$ on $Q$ (‘generalized’ has already been used for the RIPr whenever it is not attained by an element of $\mathcal{C}$ [9]). However, the density of the measure $\hat{Q}$ is only unique as an element of $M(\Omega,\mathbb{R}_{>0})$, since convergence of the densities holds in $m_{P}$. In the current work this causes no ambiguity, so that we simply refer to it as ‘the’ universal RIPr.
Note that Theorem 3 (1) implies that if there exists a $Q\in\mathcal{C}$ with $D(P\|Q\rightsquigarrow\mathcal{C})=0$, then $Q$ is the universal RIPr of $P$ on $\mathcal{C}$. This matches with the intuition that the maximum gain we can get from switching away from the ‘best’ code in $\mathcal{C}$ should be equal to zero. The following result establishes this more formally, i.e. whenever $\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})<\infty$, it must actually be equal to zero.
Proposition 2.
The following conditions are equivalent:
  1. 1.
    There exists a measure $P^{\prime}$ such that $D(P\|P^{\prime}\rightsquigarrow\mathcal{C})$ is finite.
  2. 2.
    There exists a measure $Q$ in $\mathcal{C}$ such that $D(P\|Q\rightsquigarrow\mathcal{C})$ is finite.
  3. 3.
    There exists a sequence of measures $Q_{n}\in\mathcal{C}$ such that $D(P\|Q_{n}\rightsquigarrow\mathcal{C})\to 0$ for $n\to\infty$.
To show that the universal reverse information projection exists, it is therefore enough to prove that one of these equivalent conditions hold. Which condition is easiest to check will depend on the specific setting. We now provide one example.
Proposition 3.
Assume that $\mathcal{C}$ is a convex set of probability measures that has finite minimax regret and with normalized maximum likelihood distribution $P^{\prime}\in\mathcal{C}$. Then for any probability measure $P$ that is absolutely continuous with respect to $P^{\prime}$, it holds that $D(P\|P^{\prime}\rightsquigarrow\mathcal{C})<\infty$.
One-dimensional exponential families with finite minimax regret have been classified in [18]. Proposition 3 implies that information projections exists whenever $\mathcal{C}$ is the convex hull of finitely many distributions.
Example 1.
Let $\mathcal{C}$ be a singleton whose single element $Q$ is given by the standard Gaussian and let $P$ be the standard Cauchy distribution. Since the Cauchy distribution is more heavy tailed than the Gaussian, we have that $D(P\|\mathcal{C})=\infty$. However, since both distributions have full support, it follows that
$D(P\|\mathcal{C}\rightsquigarrow Q)=D(P\|Q\rightsquigarrow Q)=0.$
By Theorem 3 (1), $Q$ is therefore the universal reverse information projection of $P$ on $\mathcal{C}$.
This example can be extended to composite $\mathcal{C}$ by considering all mixtures of the Gaussian distributions $\mathcal{N}(-1,1)$ and $\mathcal{N}(1,1)$ with mean ±$\pm 1$ and variance $1$. Proposition 3 guarantees the existence of a universal reversed information projection although the information divergence is still infinite because a Cauchy distribution is more heavy tailed than any finite mixture of Gaussian distributions. Symmetry implies that the universal reversed information projection must be equal to the uniform mixture of $\mathcal{N}(-1,1)$ and $\mathcal{N}(1,1)$.
III-A Strict sub-probability measure
We return now to the familiar setting where $P$ is a probability distribution and $\mathcal{C}$ a set of probability distributions (that is convex and closed under set-wise convergence). It is easy to verify that the RIPr $\hat{Q}$ of $P$ on $\mathcal{C}$ is then a sub-probability measure. This follows because we know that there exists a sequence $(Q_{n})_{n\in\mathbb{N}}$ in $\mathcal{C}$ such that $q_{n}$ converges point-wise $P$-a.s. to $\hat{q}$ and Fatou’s Lemma tells us
$\int_{\Omega}\hat{q}\,\mathrm{d}\mu=\int_{\Omega}\liminf_{n\to\infty}q_{n}\,\mathrm{d}\mu\leq\liminf_{n\to\infty}\int_{\Omega}q_{n}\,\mathrm{d}\mu=1.$
It is not clear a priori whether this can ever be a strict inequality. For example, if the sample space is finite, the set of probability measures is compact, so the limit of any sequence of probability measures (i.e. the reverse information projection) will also be a probability measure. The following example illustrates that this is not always the case for infinite sample spaces, and it can in fact already go wrong for a countable sample space with $D(P\|\mathcal{C})<\infty$.
Example 2.
Let $\Omega=\mathbb{N}$ and $\mathcal{F}=2^{\mathbb{N}}$. Furthermore, let $P$ denote the probability measure $\delta_{1}$ concentrated in the point $i=1$ and $\mathcal{C}$ the set of distributions $Q$ satisfying
$\sum_{i=1}^{\infty}\frac{1}{i}q(i)=\frac{1}{2}.$
This set is defined by a linear constraint, so that $\mathcal{C}$ is convex, and for any $Q\in\mathcal{C}$, we have
$q(1)+\sum_{i=2}^{\infty}\frac{1}{i}q(i)=\sum_{i=1}^{\infty}\frac{1}{i}q(i)=\frac{1}{2},$
implying that $q(1)\leq\nicefrac{{1}}{{2}}$. It follows that $D(P\|Q)=-\ln(q(1))\geq\ln(2)$. The sequence $Q_{n}=\frac{n-2}{2n-2}\delta_{1}+\frac{n}{2n-2}\delta_{n}$ satisfies $Q_{n}\in\mathcal{C}$ and
$D(P\|Q_{n})=\ln\frac{2n-2}{n-2}\to\ln(2).$
Consequently, it must hold that $D(P\|\mathcal{C})=\ln(2)$. The sequence $Q_{n}$ converges to the strict sub-probability measure $\nicefrac{{1}}{{2}}\delta_{1}$, which must therefore be the (universal) RIPr of $P$ on $\mathcal{C}$.
A more general example, which can be seen as a template to create such situations, is given in Appendix C. The common theme is that $\mathcal{C}$ is defined using only constraints of the form $\sum g(i)q(i)=c$, where $g$ is a positive function such that $\lim_{n\to\infty}g(n)=0$. The following Theorem shows that if we add a ‘dominating’ restriction to $\mathcal{C}$ such constraints cannot be violated.
Theorem 4.
Take $\Omega=\mathbb{N},\mathcal{F}=2^{\mathbb{N}}$, and let $\mathcal{C}$ be a convex set of probability distributions. Suppose that for $f_{0},f_{1}:\mathbb{N}\to\mathbb{R}_{>0}$, we have that $\sum_{i}f_{0}(i)q(i)\leq\lambda_{0}$ and $\sum_{i}f_{1}(i)q(i)=\lambda_{1}$ for all $Q\in\mathcal{C}$. If $Q_{n}$ denotes a sequence of measures in $\mathcal{C}$ that converges point-wise to some distribution $Q^{*}$, and $f_{0}$ dominates $f_{1}$ in the sense that
$\lim_{i\to\infty}\frac{f_{1}\left(i\right)}{f_{0}\left(i\right)}=0,$ (4)
then
$\sum_{i}f_{1}\left(i\right)\cdot q^{*}\left(i\right)=\lambda_{1}.$ (5)
IV Optimal E-statistics
In this section, we assume that $P$ and all $Q\in\mathcal{C}$ are probability distributions, and we are interested in the hypothesis test with $P$ as alternative and $\mathcal{C}$ as null. To this end, Theorem 3 shows that — if it exists — the likelihood ratio of $P$ and its universal RIPr is an e-statistic. A natural question is whether the optimality of the universal RIPr in terms of describing data distributed according to $P$ carries over to some sort of optimality of the e-statistic, similar as for the GRO criterion in the case that $D(P\|\mathcal{C})$. It turns out that this is true in terms of a very intuitive extension of the GRO criterion. Completely analogous to the coding story, we simply have to change from absolute to pairwise comparisons.
Definition 2.
For e-statistics $E,E^{\prime}\in\mathcal{E}_{\mathcal{C}}$, we say that $E$ is stronger than $E^{\prime}$ if the following integral is well-defined and non-negative, possibly infinite:
$\int_{\Omega}\ln\left(\frac{E}{E^{\prime}}\right)\,\mathrm{d}P,$ (6)
where we adhere to the conventions $\ln(\nicefrac{{0}}{{c}})=-\infty$ and $\ln(\nicefrac{{c}}{{0}})=\infty$ for all $c\in\mathbb{R}_{>0}$. Furthermore, an e-statistic $E^{*}\in\mathcal{E}_{\mathcal{C}}$ is the strongest e-statistic if it is stronger than any other e-statistic $E\in\mathcal{E}_{\mathcal{C}}$.
Since we assumed that there exists a $Q^{*}\in\mathcal{C}$ such that $P\ll Q^{*}$, it follows that for any e-statistic $E$ we must have $P(E=\infty)=0$, which simplifies any subsequent analyses greatly. The optimality criterion in Definition 2 can be seen as a generalization of GRO, because if $\int_{\Omega}\ln E\,\mathrm{d}P$ and $\int_{\Omega}\ln E^{\prime}\,\mathrm{d}P$ are both finite, (6) can be written as the difference between the two logarithms, thus recovering the original GRO criterion. Analogous to that case, we prove that whenever the universal RIPr exists, it leads to an optimal e-statistic.
Theorem 5.
Suppose that both $P$ and all $Q\in\mathcal{C}$ are probability distributions such that $\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})<\infty$. If $\hat{Q}$ denotes the universal RIPr of $P$ on $\mathcal{C}$, then $\hat{E}=\nicefrac{{\mathrm{d}P}}{{\mathrm{d}\hat{Q}}}$ is the strongest e-statistic. Furthermore, for any other strongest e-statistic $\tilde{E}$ we must have that $\tilde{E}=\hat{E}$ holds $P$-a.s.
The above notion of optimality comes down to the simple idea that if one e-statistic $E$ is stronger than another e-statistic $E^{\prime}$, then a test based on $E$ is more powerful than a test based on $E^{\prime}$ in the sense that there is a higher probability of rejecting a false null-hypothesis. We will formulate an asymptotic version of this idea. Suppose that we conduct the same experiment $n$ times independently to test the veracity of the hypothesis $\mathcal{C}$, resulting in outcomes $\omega_{1},\dots,\omega_{n}$. For two e-statistics $E,E^{\prime}\in\mathcal{E}_{\mathcal{C}}$, the law of large numbers states that if $P$ is true, it will almost surely hold that
$\frac{\prod_{i=1}^{n}E(\omega_{i})}{\prod_{i=1}^{n}E^{\prime}(\omega_{i})}=\exp\left(n\int_{\Omega}\ln\left(\frac{E}{E^{\prime}}\right)\,\mathrm{d}P+o(n)\right).$
It follows that if the integral $\int_{\Omega}\ln\left(\frac{E}{E^{\prime}}\right)\,\mathrm{d}P$ is positive then with high probability $E$ will give more evidence against $\mathcal{C}$ than $E^{\prime}$ if the alternative is true, i.e. a test based on $E$ will asymptotically have more power than a test based on $E^{\prime}$. We now return to Example 1, where the GRO criterion is not able to distinguish between e-variables, but we are able to do so with Definition 2 and Theorem 5.
Example 1 (continued).
In the case that $P$ is the standard Cauchy and $\mathcal{C}=\{Q\}$, where $Q$ is the standard Gaussian, it is straightforward to see that the likelihood ratio between $P$ and $Q$ is an e-statistic, i.e.
$\int_{\Omega}\frac{\mathrm{d}{P}}{\mathrm{d}Q}\,\mathrm{d}Q=\int_{\Omega}\,\mathrm{d}P=1.$
However, for the growth rate it holds that
$\int_{\Omega}\ln\left(\frac{\mathrm{d}{P}}{\mathrm{d}Q}\right)\,\mathrm{d}P=D(P\|Q)=\infty.$
The same argument can be used to show that for any $0<c\leq 1$, we have an e-statistic given by $c{\mathrm{d}P}/{\mathrm{d}Q}$, which still has infinite growth rate. The GRO criterion in Definition 1 is not able to tell which of these e-statistics is preferable. However, since $Q$ is the universal RIPr of $P$ on $\mathcal{C}$, it follows from Theorem 5 that $\nicefrac{{\mathrm{d}P}}{{\mathrm{d}Q}}$ is the strongest e-statistic, and in particular stronger than $c{\mathrm{d}P}/{\mathrm{d}Q}$ for all $10<c<$ 1.
IV-A Approximation
So far, we have discussed the existence and properties of the universal reverse information projection of $P$ on $\mathcal{C}$. However, there will be many situations where it is infeasible to compute this exact projection, as it requires solving a complex minimization problem. For example, if $\mathcal{C}$ is given by the convex hull of some parameterized family of distributions, then the universal reverse information projection might be an arbitrary mixture of elements of this family, and the minimization problem need not be convex in the parameters of the family. It will therefore often be useful to resort to algorithms that provide an approximation. Existing algorithms focus on the case that $D(P\|\mathcal{C})<\infty$, such as those proposed by Li [6, Theorem 3.3] or Csiszár and Tusnády [19, Theorem 5].
However, the convergence guarantees that these algorithms provide are in terms of information divergence, i.e. if $\hat{Q}_{i}$ is the approximation of the projection after $i$ iterations, then $D(P\|\hat{Q}_{i})\to D(P\|\mathcal{C})$. If we want to use such an approximation for hypothesis testing, this is not enough: we need that $\nicefrac{{p}}{{q_{i}}}$ gets closer and closer to being an e-statistic. The following theorem gives a condition under which this is true.
Proposition 4.
Assume that $D(P\|\mathcal{C})<\infty$, fix $Q,Q^{\prime}\in\mathcal{C}$, and let $C_{q,q^{\prime}}>0$ be such that
$\int_{\Omega}\left(\frac{q^{\prime}}{q}\right)^{2}\,\mathrm{d}P=C_{q,q^{\prime}}.$ (7)
Then if we set $\delta:=D(P\|Q\rightsquigarrow\mathcal{C})$, we have
$\int_{\Omega}\frac{p}{q}\,\mathrm{d}Q^{\prime}=\int_{\Omega}\frac{q^{\prime}}{q}\,\mathrm{d}P\leq 1+4\max\left\{\left(C_{q,q^{\prime}}\delta\right)^{\nicefrac{{1}}{{2}}},\delta\right\}.$ (8)
Moreover, if, for some $K\geq 0$, $D(P\|Q^{\prime}\rightsquigarrow\mathcal{C})\leq K\delta$ and there exists a constant $c_{1}>0$ such that either $\nicefrac{{q^{\prime}}}{{q}}\leq c_{1}$ $P$-almost surely or $\nicefrac{{q}}{{q^{\prime}}}\leq c_{1}$ $P$-almost surely, then
$\int_{\Omega}\frac{p}{q}\,\mathrm{d}Q^{\prime}=\int_{\Omega}\frac{q^{\prime}}{q}\,\mathrm{d}P\leq 1+(4\left(2c_{1}\max\{1,K\}\right)^{{}^{\nicefrac{{1}}{{2}}}}+2)\delta.$ (9)
We conjecture that a similar result will be true in the more general setting when $\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})<\infty$ even when $D(P\|\mathcal{C})=\infty$. Furthermore, while the conditions (8) and (9) may seem very strong, the following example shows that in general such a condition is needed for the convergence to hold.
Example 3.
Let $\mathcal{Q}$ represent the family of geometric distributions on $\Omega=\mathbb{N}_{0}$ and let $\mathcal{C}=\text{conv}(\mathcal{Q})$. The elements of $\mathcal{Q}$ are denoted by $Q_{\theta}$ with density $q_{\theta}(n)=\theta^{n}(1-\theta)$, where $\theta\in[0,1)$ denotes the probability of failure. For simplicity, assume that $P\in\mathcal{Q}$ so that the reverse information projection of $P$ on $\mathcal{C}$ is equal to $P$. Take for example $P=Q_{\nicefrac{{1}}{{2}}}$, then for any $\theta,\theta^{\prime}\in[0,1)$
$\int_{\Omega}\frac{q_{\theta^{\prime}}}{q_{\theta}}\,\mathrm{d}P$ $=\sum_{n=0}^{\infty}\left(\frac{1}{2}\frac{\theta^{\prime}}{\theta}\right)^{n}\frac{1}{2}\frac{1-\theta^{\prime}}{1-\theta}$ (10)
=$=\begin{cases}\frac{1}{1-\frac{\theta^{\prime}}{2\theta}}\cdot\frac{1}{2}\frac{1-\theta^{\prime}}{1-\theta},&\text{\ if\ }\theta^{\prime}<2\theta;\\ \infty,&\text{\ otherwise;}\end{cases}$ (11)
whereas
$D(P\|Q_{\theta})$ $=\sum_{n=0}^{\infty}\left(\frac{1}{2}\right)^{n+1}\left(-n\log(2\theta)-\log 2(1-\theta)\right)$
$=\log\frac{\nicefrac{{1}}{{2}}}{1-\theta}+\log\frac{\nicefrac{{1}}{{2}}}{\theta}.$
Now consider a sequence $\nicefrac{{1}}{{3}}<\theta_{1}<\theta_{2}<\theta_{3}\ldots$ that converges to $\nicefrac{{1}}{{2}}$. Then by the above,
$D(P\|Q_{\theta_{i}})\to 0=D(P\|\mathcal{C}).$
We also see that for all $i$ and all $\theta^{\prime}\in[2\theta_{i},1)$, we have
$\int_{\Omega}q_{\theta^{\prime}}/q_{\theta_{i}}\mathrm{d}P=\infty,$
i.e. for all $i$ we have $\sup_{\theta^{\prime}\in[0,1)}\int_{\Omega}q_{\theta^{\prime}}/q_{\theta_{i}}\mathrm{d}P=\infty$. This shows that in general, a condition such as (7) is necessary.
V Future and Related Work
The results presented thus far suggest various avenues for further research of which we now discuss two. First, Theorem 3 is formulated for general measures so one may ask for an interpretation of the universal RIPr in the case that $P$ and $\mathcal{C}$ are not probability measures. If $\Omega$ is finite and $\lambda$ is a measure on $\Omega$, then we may define a probability measure $Po(\lambda)$ as the product measure $Po(\lambda)=\bigotimes_{\omega\in\Omega}Po\left(\lambda(\omega)\right)$, where $Po\left(\lambda(\omega)\right)$ denotes the Poisson distribution with mean $\lambda(\omega)$. With this definition we get
$D(P\|Q\rightsquigarrow Q^{\prime})=D(Po(P)\|Po(Q)\rightsquigarrow Po(Q^{\prime})).$
Furthermore, it can be shown that if the universal RIPr $\hat{Q}$ of $P$ on $\mathcal{C}$ exists and is an element of $\mathcal{C}$, then $Po(\hat{Q})$ is also the universal RIPr of $Po(P)$ on the convex hull of $\mathcal{C}^{\prime}:=\{Po(Q)|Q\in\mathcal{C}\}$. Consequently, $\nicefrac{{Po(P)}}{{Po(\hat{Q})}}$ can be thought of as an e-statistic for $\mathcal{C}^{\prime}$. More work is needed to determine whether this interpretation has any applications and if it can be generalized to arbitrary $\Omega$.
Second, even if $D(P\|\mathcal{C})=\infty$, the Rényi divergence $D_{\alpha}\left(P\|Q\right)$ (see e.g. [20]) may be a well-defined non-negative real number for $\alpha\in(0,1)$ and $Q\in\mathcal{C}$. These Rényi divergences are jointly convex in $P$ and $Q$ [20] and for each $10<\alpha<$ 1 one may define a reversed Rényi projection $\hat{Q}_{\alpha}$ of $P$ on $\mathcal{C}$ [21]. If it exists, it can be shown that this distribution will satisfy
$\int_{\Omega}\left(\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}_{\alpha}}\right)^{\alpha}\,\mathrm{d}Q\leq 1$
for all $Q\in\mathcal{C}$, i.e. $\left(\nicefrac{{\mathrm{d}P}}{{\mathrm{d}\hat{Q}_{\alpha}}}\right)^{\alpha}$ is an e-statistic. We conjecture that the projections $\hat{Q}_{\alpha}$ will converge to the universal RIPr for $\alpha$ tending to 1, which might lead to further applications.
Except for the present paragraphs, this article is near-identical to the paper that we submitted to ISIT 2023 under the same title. After acceptance of that paper, we discovered that there is some overlap between work concurrently done by Zhang et al. [22] and our results on the existence of optimal e-statistics. In particular, they show that if $\mathcal{C}$ is a convex polytope, then there exists an e-statistic in the form of a likelihood ratio between two unspecified measures. Since a convex polytope contains the uniform mixture of its vertices, which can be shown to have finite information gain, this also follows from our Theorem 3. They furthermore show that if the alternative is also a convex polytope $\mathcal{A}$, then at least one of their e-statistics in the form of a likelihood ratio satisfies $\inf_{P\in\mathcal{A}}\int_{\Omega}\ln E\,\mathrm{d}P>0$. More work is needed to determine whether our construction leads to a similar result for composite alternatives. Finally, it should be noted that the techniques used to prove their results appear to be of a completely different nature than the ones used in this paper, as they rely mostly on classical results in convex geometry together with results on optimal transport.
Appendix A Proofs
A-A Proof of Proposition 1
Proposition 1 is a direct corollary of two auxiliary results, which we state here.
Lemma 1.
For $x,y\in\mathbb{R}_{>0}$, we have
$\left|\ln(x)-\ln(y)\right|=g(m_{\gamma}^{2}(x,y)),$
where $g$ denotes the function
$g(t)=2t+2\ln\left(1+\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}\right).$
The function $g$ is concave and satisfies $g(t)\geq 2t$.
Proof.
Let $m=\frac{x+y}{2}$. Our goal is to determine the function $g$ function such that
$\left|\ln(x)-\ln(y)\right|=g(m^{2}_{\gamma}(x,y)).$
We first rewrite the right-hand side
$g(m^{2}_{\gamma}(x,y))$ $=g\left(\ln\left(m\right)-\frac{1}{2}\ln\left(x\right)-\frac{1}{2}\ln\left(y\right)\right)$
$=g\left(\frac{1}{2}\ln\left(\frac{m^{2}}{x\cdot y}\right)\right)$
$=g\left(\frac{1}{2}\ln\left(\frac{\left(\frac{m}{y}\right)^{2}}{\frac{x}{y}}\right)\right)$
$=g\left(\frac{1}{2}\ln\left(\frac{\left(\frac{1+\frac{x}{y}}{2}\right)^{2}}{\frac{x}{y}}\right)\right).$
Plugging this back in and replacing $\frac{x}{y}$ by $w$ leads to
$\left|\ln\left(w\right)\right|=g\left(\frac{1}{2}\ln\left(\frac{\left(\frac{1+w}{2}\right)^{2}}{w}\right)\right)$
Then we solve the equation
$\frac{1}{2}\ln\left(\frac{\left(\frac{1+w}{2}\right)^{2}}{w}\right)=t,$
which gives
$w=2\exp\left(2t\right)-1+2\cdot\left(\exp\left(4t\right)-\exp\left(2t\right)\right)^{\nicefrac{{1}}{{2}}}$
$g\left(t\right)$ $=\ln\left(2\exp\left(2t\right)-1+2\cdot\left(\exp\left(4t\right)-\exp\left(2t\right)\right)^{\nicefrac{{1}}{{2}}}\right)$
$=2t+\ln\left(2-\exp\left(-2t\right)+2\cdot\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}\right)$
$=2t+2\ln\left(1+\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}\right).$
The derivatives of $g$ are
$g^{\prime}(t)$ $=2+2\frac{\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{-1}}{{2}}}\exp\left(-2t\right)}{1+\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}}$
$g^{\prime\prime}(t)$ $=\frac{-\exp\left(-\nicefrac{{t}}{{2}}\right)}{2^{\nicefrac{{1}}{{2}}}\left(\sinh{t}\right)^{\nicefrac{{3}}{{2}}}}.$
We see that $g^{\prime\prime}(t)<0$ and conclude that $g$ is concave.
We have
$g\left(t\right)$ $=2t+\ln\left(2-\exp\left(-2t\right)+2\cdot\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}\right)$
$=2t+\ln\left(1+\left(1-\exp\left(-2t\right)\right)+2\cdot\left(1-\exp\left(-2t\right)\right)^{\nicefrac{{1}}{{2}}}\right)$
$\geq 2t$
because $1-\exp\left(-2t\right)\geq 0$. ∎
Lemma 2.
Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of elements of $\mathcal{M}(\Omega,\mathbb{R}_{>0})$, then
$\limsup_{m,n\to\infty}m_{P}(f_{m},f_{n})=0\Rightarrow\limsup_{m,n\to\infty}\int_{\Omega}\left|\ln\left(\frac{f_{m}}{f_{n}}\right)\right|\,\mathrm{d}P=0.$
Proof.
By Lemma 1, we have for $m,n\in\mathbb{N}$,
$\int_{\Omega}\left|\ln\left(\frac{f_{n}}{f_{m}}\right)\right|\,\mathrm{d}P$ $=\int_{\Omega}g(m^{2}_{\gamma}(f_{n},f_{m}))\,\mathrm{d}P$
$\leq g\left(\int_{\Omega}m^{2}_{\gamma}(f_{n},f_{m})\,\mathrm{d}P\right)$
$=g\left(m^{2}_{P}\left(f_{n},f_{m}\right)\right).$
The result follows by continuity of $g$. ∎
A-B Proofs for Section III
Before giving the intended results, we note that we introduced $m_{P}$ as the averaged Bregman divergence associated with $\gamma(x)=x-1-\ln(x)$. For the proof, it will be useful to also define the Bregman divergence associated with $\gamma(x)=x-1-\ln(x)$ itself, which is the so-called Itakura-Saito divergence. For $f,g\in\mathcal{M}\left(\Omega,(0,\infty)\right)$, it is given by
$IS_{P}(f,g)=\int_{\Omega}\left(\frac{f}{g}-1-\ln\frac{f}{g}\right)\,\mathrm{d}P.$
By definition, it holds that
$m_{P}^{2}(f,g)=\frac{1}{2}IS\left(f,\frac{f+g}{2}\right)+\frac{1}{2}IS\left(g,\frac{f+g}{2}\right).$
Furthermore, for $Q\in\mathcal{C}$, we have $IS_{P}(q,p)=D(P\|Q)$. We now state an auxiliary result before giving the proofs for Section III.
Lemma 3.
For $Q_{1},Q_{2}\in\mathcal{C}$, we have
$m_{P}^{2}(q_{1},q_{2})\leq\frac{D(P\|Q_{1}\rightsquigarrow\mathcal{C})+D(P\|Q_{2}\rightsquigarrow\mathcal{C})}{2}.$
Proof.
Let $\bar{Q}$ denote the midpoint between $Q_{1}$ and $Q_{2}$. Then we have
$\frac{D(P\|Q_{1}\rightsquigarrow\mathcal{C})+D(P\|Q_{2}\rightsquigarrow\mathcal{C})}{2}\\ =\frac{\sup_{Q\in\mathcal{C}}D(P\|Q_{1}\rightsquigarrow Q)+\sup_{Q\in\mathcal{C}}D(P\|Q_{2}\rightsquigarrow Q)}{2}\\ \geq\frac{D(P\|Q_{1}\rightsquigarrow\bar{Q})+D(P\|Q_{2}\rightsquigarrow\bar{Q})}{2}=m_{P}^{2}(q_{1},q_{2}).$
We deviate slightly from the order of the results in Section III and first state the proof of Proposition 2, so that we can use its results in the proof of Theorem 3.
Proof of Proposition 2.
The implications $(1)\to(2)\to(3)$ are obvious. Assume that $P^{\prime}$ is a measure such that $-\infty<D(P\|P^{\prime}\rightsquigarrow\mathcal{C})<\infty$. Then there exists a sequence of measures $Q_{n}\in\mathcal{C}$ such that
$D(P\|P^{\prime}\rightsquigarrow Q_{n})\to D(P\|P^{\prime}\rightsquigarrow\mathcal{C})$
for $n\to\infty$. Without loss of generality we may assume that $-\infty<D(P\|P^{\prime}\rightsquigarrow Q_{n})<\infty$ for all $n$. The result follows because
$D(P\|P^{\prime}\rightsquigarrow\mathcal{C})=D(P\|P^{\prime}\rightsquigarrow Q_{n})+D(P\|Q_{n}\rightsquigarrow\mathcal{C})$
and all involved quantities are finite. ∎
Proof of Theorem 3 1.
Let $(Q_{n})_{n\in\mathcal{C}}$ denote a sequence in $\mathcal{C}$ such that
$\lim_{n\to\infty}D(P\|Q_{n}\rightsquigarrow\mathcal{C})=\inf_{Q\in\mathcal{C}}D(P\|Q\rightsquigarrow\mathcal{C})=0,$
where the last equality follows from Proposition 2. By Lemma 3, we have for $m,n\in\mathbb{N}$ that
$m_{P}^{2}(q_{m},q_{n})\leq\\ \frac{D(P\|Q_{m}\rightsquigarrow\mathcal{C})+D(P\|Q_{n}\rightsquigarrow\mathcal{C})}{2}.$ (12)
It follows that $(q_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence with respect to $m_{P}$, so that $(q_{n})_{n\in\mathbb{N}}$ converges to some function $\hat{q}$ in $m_{P}$. The latter follows from the completeness of $(\mathcal{M}\left(\Omega,(0,\infty)\right),m_{P})$, i.e. Proposition 1.
Furthermore, suppose that $(Q^{\prime}_{n})_{n\in\mathcal{C}}$ is another sequence in $\mathcal{C}$ such that
$\lim_{n\to\infty}D(P\|Q^{\prime}_{n}\rightsquigarrow\mathcal{C})=0.$
Then $Q_{1},Q^{\prime}_{1},Q_{2},Q^{\prime}_{2},Q_{3},Q^{\prime}_{3},\dots$ is also a Cauchy sequence that converges and since a Cauchy sequence can only converge to a single element this implies the desired uniqueness. ∎
Proof of Theorem 3 2.
The equality
$\int_{\Omega}\ln\frac{p^{\prime}}{\hat{q}}\,\mathrm{d}P=\lim_{n\to\infty}\int_{\Omega}\ln\frac{p^{\prime}}{q_{n}}\,\mathrm{d}P$
follows from Theorem 3 1 together with the fact that convergence of $q_{n}$ in $m_{P}$ implies convergence of the logarithms in $L_{1}(P)$. ∎
Proof of Theorem 3 3.
Let $(Q_{n})_{n\in\mathcal{C}}$ denote a sequence in $\mathcal{C}$ such that
$\lim_{n\to\infty}D(P\|Q_{n}\rightsquigarrow\mathcal{C})=0.$
Without loss of generality, we may assume that $D(P\|Q_{n}\rightsquigarrow\mathcal{C})<\infty$ for all $n$. Further, $q_{n}$ converges to $\hat{q}$ point-wise $P$-almost surely (convergence in $m_{P}$ implies convergence of the logarithms in $L_{1}(P)$, which gives the existence of an almost surely converging sub-sequence).
Let $\tilde{Q}=(1-t)Q_{1}+tQ$ for fixed $Q\in\mathcal{C}$ and fixed $10<t<$ 1. Let $Q_{n,s}$ denote the convex combination $Q_{n,s}=(1-s_{n})Q_{n}+s_{n}\tilde{Q}$ and $s_{n}\in[0,1]$. By Theorem 3 1, we know that there exists some $\hat{Q}$ such that $q_{n}\to\hat{q}$ in $m_{P}$.
Since $Q_{n,s}\in\mathcal{C}$ by convexity, we have that $D(P\|Q_{n}\rightsquigarrow Q_{n,s})\leq D(P\|Q_{n}\rightsquigarrow\mathcal{C})$. We get
$D(P\|Q_{n}\rightsquigarrow Q_{n,s})=\\ s_{n}D(P\|Q_{n}\rightsquigarrow\tilde{Q})+s_{n}IS_{P}\left(\tilde{q},q_{n,s}\right)+(1-s_{n})IS_{P}\left(q_{n},q_{n,s}\right)\\ \geq s_{n}D(P\|Q_{n}\rightsquigarrow\tilde{Q})+s_{n}IS_{P}(\tilde{q},q_{n,s}).$
Hence
$s_{n}D(P\|Q_{n}\rightsquigarrow\tilde{Q})+s_{n}IS_{P}(\tilde{q},q_{n,s})\leq D(P\|Q_{n}\rightsquigarrow\mathcal{C}).$
Division by $s_{n}$ gives
$D(P\|Q_{n}\rightsquigarrow\tilde{Q})+IS_{P}(\tilde{q}q_{n,s})\leq\frac{D(P\|Q_{n}\rightsquigarrow\mathcal{C})}{s_{n}}.$
Choosing $s_{n}=D(P\|Q_{n}\rightsquigarrow\mathcal{C})^{\nicefrac{{1}}{{2}}}$, this gives
$D(P\|Q_{n}\rightsquigarrow\tilde{Q})+IS_{P}(\tilde{q},q_{n,s})$ $\leq s_{n}^{\nicefrac{{1}}{{2}}}.$
Then we get
$IS_{P}(\tilde{q},q_{n,s})$ $\leq D(P\|\tilde{Q}\rightsquigarrow Q_{n})+s_{n}^{\nicefrac{{1}}{{2}}}.$
$\int_{\Omega}\left(\frac{\tilde{q}}{q_{n,s}}+\ln\frac{q_{n,s}}{q_{n}}\right)\,\mathrm{d}P$ $\leq P(\Omega)+\tilde{Q}(\Omega)-Q_{n}(\Omega)+s_{n}^{\nicefrac{{1}}{{2}}}.$
Writing $q_{n}$ as $\frac{q_{n,s}-s_{n}\tilde{q}}{1-s}$, we see
$\ln\frac{q_{n,s}}{q_{n}}$ $=\ln\frac{q_{n,s}}{\frac{q_{n,s}-s_{n}\tilde{q}}{1-s}}$
$=\ln(1-s_{n})-\ln\frac{q_{n,s}-s_{n}\tilde{q}}{q_{n,s}}$
$=\ln(1-s_{n})-\ln\left(1-s_{n}\frac{\tilde{q}}{q_{n,s}}\right)$
$\geq\ln(1-s_{n})+s_{n}\frac{\tilde{q}}{q_{n,s}}.$
Hence
$\ln(1-s_{n})P(\Omega)+(1+s_{n})\int_{\Omega}\frac{\tilde{q}}{q_{n,s}}\,\mathrm{d}P\\ \leq P(\Omega)+\tilde{Q}(\Omega)-Q_{n}(\Omega)+s_{n}^{\nicefrac{{1}}{{2}}}.$
As $\lim_{n\to\infty}s_{n}=0$, taking the limit inferior as $n\to\infty$ on both sides gives
$\liminf_{n\to\infty}\int_{\Omega}\frac{\tilde{q}}{q_{n,s}}\,\mathrm{d}P\leq P(\Omega)+\tilde{Q}(\Omega)-\liminf_{n\to\infty}Q_{n}(\Omega).$
An application of Fatou’s lemma gives
$\int_{\Omega}\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}Q\leq P(\Omega)+\tilde{Q}(\Omega)-\liminf_{n\to\infty}Q_{n}(\Omega).$
Since $\tilde{Q}=(1-t)Q_{1}+tQ$ get the inequality
$\int_{\Omega}$ $\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}\left(\left(1-t\right)Q+tQ_{m}\right)$
$\leq P(\Omega)+(1-t)Q_{1}(\Omega)+tQ(\Omega)-\liminf_{n\to\infty}Q_{n}(\Omega),$
$\left(1-t\right)$ $\int_{\Omega}\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}Q+t\int_{\Omega}\frac{\mathrm{d}P}{\mathrm{d}\hat{Q}}\,\mathrm{d}Q_{m}$
$\leq P(\Omega)++(1-t)Q_{1}(\Omega)+tQ(\Omega)-\liminf_{n\to\infty}Q_{n}(\Omega).$
Finally we let $t$ tend to zero and obtain the desired result. ∎
Proof of Proposition 3.
Since $P^{\prime}$ is the normalized maximum likelihood distribution we have $\sup_{Q}\sup_{\omega}\ln{\frac{\mathrm{d}Q}{\mathrm{d}P^{\prime}}}<\infty.$ In particular
$\sup_{Q\in\mathcal{C}}D(P\|P^{\prime}\rightsquigarrow Q)$ $=\sup_{Q\in\mathcal{C}}\int_{\Omega}\ln{\frac{\mathrm{d}Q}{\mathrm{d}P^{\prime}}}\,\mathrm{d}P$
$\leq\sup_{Q\in\mathcal{C}}\sup_{\omega}\ln{\frac{\mathrm{d}Q}{\mathrm{d}P^{\prime}}(\omega)}$
<<$\infty$.< ∞ .
A-C Proof of Theorem 4
Proof of Theorem 4.
For any $a\in\mathbb{R}$ we have
$f_{0}\left(i\right)+a\cdot f_{1}\left(i\right)=f_{0}\left(i\right)\cdot\left(1+a\cdot\frac{f_{1}\left(i\right)}{f_{0}\left(i\right)}\right).$ (13)
Since $\frac{f_{1}\left(i\right)}{f_{0}\left(i\right)}\to 0$ for $i\to\infty$ we have that $f_{0}\left(i\right)+a\cdot f_{1}\left(i\right)\geq 0$ for $i$ sufficiently large. Therefore, we can apply Fatou’s lemma to the function and obtain
$\sum f_{0}\left(i\right)\cdot q^{*}\left(i\right)+a\cdot\sum f_{1}\left(i\right)\cdot q^{*}\left(i\right)\\ =\sum\left(f_{0}\left(i\right)+a\cdot f_{1}\left(i\right)\right)\cdot q^{*}\left(i\right)\\ =\sum\liminf_{n\to\infty}\left(f_{0}\left(i\right)+a\cdot f_{1}\left(i\right)\right)\cdot q_{n}\left(i\right)\\ \leq\liminf_{n\to\infty}\sum_{i}\left(f_{0}\left(i\right)+a\cdot f_{1}\left(i\right)\right)\cdot q_{n}\left(i\right)\\ =\liminf_{n\to\infty}\left(\sum_{i}f_{0}\left(i\right)\cdot q_{n}\left(i\right)+a\cdot\sum_{i}f_{1}\left(i\right)\cdot q_{n}\left(i\right)\right)\\ =\liminf_{n\to\infty}\left(\lambda_{0}+a\cdot\lambda_{1}\right)=\lambda_{0}+a\cdot\lambda_{1}.$
Hence
$a\cdot\left(\sum f_{1}\left(i\right)\cdot q^{*}\left(i\right)-\lambda_{1}\right)\leq\lambda_{0}-\sum f_{0}\left(i\right)\cdot q^{*}\left(i\right).$ (14)
This inequality should hold for all $a\in\mathbb{R}$, which is only possible if
$\sum f_{1}\left(i\right)\cdot q^{*}\left(i\right)-\lambda_{1}$ $=0.$
$\sum f_{1}\left(i\right)\cdot q^{*}\left(i\right)$ $=\lambda_{1}.$
Appendix B Proofs for Section IV
B-A Proof of Theorem 5
Proof of Theorem 5.
Firstly, since $\hat{E}>0$ holds $P$-almost surely, we have that $\hat{E}$ is stronger than any $E^{\prime}\in\mathcal{E}_{\mathcal{C}}$ with $P(E^{\prime}=0)>0$.
Secondly, let $E\in\mathcal{E}_{\mathcal{C}}$ be an e-statistic for which $E>0$ holds $P$-almost surely. Furthermore, let $Q_{n}$ be a sequence of measures in $\mathcal{C}$ such that $D(P\|Q_{n}\mathcal{C})\to 0$. We can define a sequence of sub-probability measures $R_{n}$ by $R_{n}(F)=\int_{F}E\,\mathrm{d}\hat{Q}_{n}$, which satisfies $\nicefrac{{\mathrm{d}R_{n}}}{{\mathrm{d}\hat{Q}_{n}}}=E$. We see
$\int_{\Omega}\ln\left(\frac{\hat{E}}{E}\right)\,\mathrm{d}P$ $=\int_{\Omega}\ln\left(\frac{\mathrm{d}Q_{n}}{\mathrm{d}\hat{Q}}\right)\,\mathrm{d}P+D(P\|R_{n})$
$+(P(\Omega)-R_{n}(\Omega))$
$\geq\int_{\Omega}\ln\left(\frac{\mathrm{d}Q_{n}}{\mathrm{d}\hat{Q}}\right)\,\mathrm{d}P.$
The last expression goes to zero as $n\to\infty$, so we see that $\hat{E}$ is stronger than $E$. ∎
B-B Proof of Proposition 4
Proof of Proposition 4.
Without loss of generality, assume that $\int_{\Omega}q^{\prime}/q\,\mathrm{d}P=1+\epsilon$ for some $\epsilon>0$. For the sake of brevity, we write $c:=C_{q,q^{\prime}}$. We now define a function $g:[0,1]\to\mathbb{R}_{\geq 0}$ as
$g(\alpha):=D(P\|(1-\alpha)Q+\alpha Q^{\prime})-D(P\|\mathcal{C}).$
Notice that $g(0)=\delta$ and $g(\alpha)\geq 0$, since $(1-\alpha)Q+\alpha Q^{\prime}\in\mathcal{C}$. This function and its derivatives will guide the rest of the proofs, and we now list some properties that we will need:
$g^{\prime}(\alpha)$ $:=\frac{\mathrm{d}}{\mathrm{d}\alpha}g(\alpha)=\int_{\Omega}\frac{q-q^{\prime}}{(1-\alpha)q+\alpha q^{\prime}}\,\mathrm{d}P,\text{\ so \ }$ (15)
$g^{\prime}(0)$ $=\int_{\Omega}\left(1-\frac{q^{\prime}}{q}\right)\,\mathrm{d}P=-\epsilon,$ (16)
$g^{\prime\prime}(\alpha)$ $:=\frac{\mathrm{d}^{2}}{\mathrm{d}\alpha^{2}}g(\alpha)=\int_{\Omega}\left(\frac{q^{\prime}-q}{(1-\alpha)q+\alpha q^{\prime}}\right)^{2}\,\mathrm{d}P,\text{\ so \ }$ (17)
$g^{\prime\prime}(0)$ $=\int_{\Omega}\left(1-\frac{q^{\prime}}{q}\right)^{2}\,\mathrm{d}P=1-2(1+\epsilon)+c\text{\ and \ }$ (18)
$0\leq g^{\prime\prime}(\alpha)$ $\leq\frac{1}{(1-\alpha)^{2}}g^{\prime\prime}(0).$ (19)
By Taylor’s theorem, we have for any $\alpha\in[0,\nicefrac{{1}}{{2}}]$ that
$g^{\prime}(\alpha)$ $\leq g^{\prime}(0)+\alpha\max_{0\leq\alpha^{\circ}\leq\alpha}g^{\prime\prime}(\alpha^{\circ})$
$\leq g^{\prime}(0)+\alpha 4g^{\prime\prime}(0)\leq-\epsilon+4\alpha c,$
where we use the properties derived above. It follows that $g^{\prime}(\alpha)\leq-\nicefrac{{\epsilon}}{{2}}$ for all $\alpha\in[0,\alpha^{*}]$ where $\alpha^{*}=\min\{\nicefrac{{1}}{{2}},\nicefrac{{\epsilon}}{{8c}}\}$. Consider the case that $\alpha^{*}=\nicefrac{{\epsilon}}{{8c}}<\nicefrac{{1}}{{2}}$. By Taylor’s theorem,
$g(\alpha^{*})\leq g(0)+\alpha^{*}\max_{\alpha^{\circ}\in[0,\alpha^{*}]}g^{\prime}(\alpha^{\circ})\leq\delta-\frac{\alpha^{*}\epsilon}{2}\leq\delta-\frac{\epsilon^{2}}{16c}.$
Since $g(\alpha^{*})\geq 0$ it follows that $\epsilon^{2}\leq\delta 16(1+c)$, so we can derive $\epsilon\leq 4\left(\delta c\right)^{\nicefrac{{1}}{{2}}}$. Analogously, if $\alpha^{*}=\nicefrac{{1}}{{2}}$, i.e. $\nicefrac{{\epsilon}}{{8(1+c)}}>\nicefrac{{1}}{{2}}$, we can derive $\epsilon\leq 4\delta$, and the result follows.
For the second part, we only consider the case that $q^{\prime}/q\leq c_{1}$ $P$-almost surely (the case with a bound on $q/q^{\prime}$ follows by completely symmetric reasoning). A simple computation gives, for all $0\leq\alpha\leq 1$, $(1-\alpha)q+\alpha q\leq c_{1}q$. From (17) and (19) we now see that
$g^{\prime\prime}(\alpha)$ $\geq\frac{g^{\prime\prime}(0)}{c_{1}}$ for all $\text{for all $0\leq\alpha\leq 1$};$ (20)
$g^{\prime\prime}(\alpha)$ $\leq\frac{g^{\prime\prime}(0)}{(1-\alpha)^{2}}$ for all $\text{for all $0\leq\alpha\leq\frac{1}{2}$}.$ (21)
Further note that for any $10<\alpha^{*}<$ 1, a Taylor approximation around $\alpha^{*}$ gives
$g(1)\geq g(\alpha^{*})+(1-\alpha^{*})g^{\prime}(\alpha^{*})+\frac{1}{2}\left(1-\alpha^{*}\right)^{2}\min_{\alpha\in[\alpha^{*},1]}g^{\prime\prime}(\alpha).$
If $g^{\prime}(\nicefrac{{1}}{{2}})\geq 0$, we apply this with $\alpha^{*}=\nicefrac{{1}}{{2}}$; using $g(\alpha^{*})\geq 0$ and (20) we find
$K\delta=g(1)\geq\frac{g^{\prime\prime}(0)}{8c_{1}}.$
If $g^{\prime}(\nicefrac{{1}}{{2}})\leq 0$, we perform analogous steps, Tayloring $g(0)$ instead of $g(1)$ around $\alpha^{*}=\nicefrac{{1}}{{2}}$, to find $\delta=g(0)\geq g^{\prime\prime}(0)c_{1}^{-1}/8$. Taken together this gives $g^{\prime\prime}(0)\leq 8c_{1}K^{\prime}\delta$, with $K^{\prime}=\max\{K,1\}$ and hence $g^{\prime\prime}(\alpha)\leq 8c_{1}K^{\prime}\delta/(1-\alpha)^{2}$ by (21), for $0\leq\alpha\leq\nicefrac{{1}}{{2}}$.
If $g^{\prime}(0)>0$ the desired result holds. So we may assume $g^{\prime}(0)=-c_{2}\delta$ for some $c_{2}>0$. Then (again by Taylor), on $\alpha\in[0,\nicefrac{{1}}{{2}}]$,
$g^{\prime}(\alpha)\leq-c_{2}\delta+\alpha\max_{\alpha^{\circ}\in[0,\alpha]}g^{\prime\prime}(\alpha^{\circ})\leq-c_{2}\delta+\frac{\alpha}{(1-\alpha)^{2}}8c_{1}K^{\prime}\delta,$
and, yet again by Taylor, and using $g(0)=\delta$ and $g(\alpha)\geq 0$ and $g^{\prime}(\alpha)$ is increasing on $\alpha\in[0,\nicefrac{{1}}{{2}}]$, we find that for $0\leq\alpha\leq\nicefrac{{1}}{{3}}$:
$0\leq g(\alpha)\leq\delta+\alpha\max_{\alpha^{\circ}\in[0,\alpha]}g^{\prime}(\alpha^{\circ})\\ \leq\delta+\alpha\left(-c_{2}\delta+\frac{\alpha}{(1-\alpha)^{2}}8c_{1}K^{\prime}\delta\right),$
which implies that $c_{2}\leq\frac{\alpha}{(1-\alpha)^{2}}8c_{1}K^{\prime}+\frac{1}{\alpha}$. Plugging in
$\alpha=1/(1+\left(8c_{1}\right)^{\nicefrac{{1}}{{2}}})$
(chosen to make both terms equal, and allowed because $c_{1}\geq 1$ and $K^{\prime}\geq 1$ so that $\alpha<\nicefrac{{1}}{{3}}$) we get $c_{2}\leq 4\left(2c_{1}K^{\prime}\right)^{\nicefrac{{1}}{{2}}}+2,$ and the desired result follows. ∎
Appendix C Supplementary Material
In this section, we discuss a general way to construct a measure $P$ and convex set of distributions $\mathcal{C}$ such that the reverse information projection of $P$ on $\mathcal{C}$ is a strict sub-probability measure. For simplicity, we take $\Omega=\mathbb{N}$ and $\mathcal{F}=2^{\mathbb{N}}$, though the idea should easily translate to more general settings.
Proposition 5.
Let $g:\mathbb{N}\to\mathbb{R}_{>0}$ be a function, and let $\mathcal{C}$ denote the set of measures $\{Q:\sum_{i}g\left(i\right)q(i)\leq\nu\}$ for some $\nu>0$. Then for any $P$ that is not in $\mathcal{C}$ we have that $E\left(i\right)=\nicefrac{{g\left(i\right)}}{{\nu}}$ is the optimal e-statistic.
Proof.
The extreme points in $\mathcal{C}$ are the measure with total mass 0 and measures of the form $\frac{\nu}{g\left(i\right)}\delta_{i}$, i.e. measures concentrated in single points. An e-statistic $E$ must satisfy
$\sum_{j}E\left(j\right)\frac{\nu}{g\left(i\right)}\delta_{i}\left(j\right)\leq 1$
or, equivalently, $E\left(i\right)\frac{\nu}{g\left(i\right)}\leq 1$. Hence $E\leq\nicefrac{{g}}{{\nu}}$ so the optimal e-statistic is $\nicefrac{{g}}{{\nu}}$. ∎
Let $g:\mathbb{N}\to\mathbb{R}_{>0}$ be any function that satisfies
$\lim_{n\to\infty}g(n)=0.$
Furthermore, let $P$ denote a probability measure on the natural numbers such that
$\sum_{i}\frac{p(i)}{g(i)}=c$
for some $c\in\mathbb{R}_{>0}$. For $\nu\in(0,1/c)$ and let $\mathcal{C}_{\nu}$ denote the set of measures $\{Q:\sum_{i}g\left(i\right)q(i)\leq\nu\}$. Note that we do not yet require all measures in $\mathcal{C}_{\nu}$ to be probability measures so that the set $\mathcal{C}_{\nu}$ is compact. It follows that there exists a unique element of $\mathcal{C}$ that minimizes $\sum_{i}p(i)\ln\left(p(i)/q(i)\right)$.
The optimal e-statistic is $E_{\nu}=\nicefrac{{g}}{{\nu}}$, and we may define the measure $\hat{Q}_{\nu}$ by
$\hat{q}_{\nu}(i)=\frac{p(i)}{E_{\nu}(i)}=\nu p(i)/g(i),$
and we can check that $\hat{Q}_{\nu}\in\mathcal{C}$. Hence $\hat{Q}_{\nu}$ minimizes $\sum_{i}p(i)\ln\left(p(i)/q(i)\right)$.
This is a strict sub-probability measure:
$\sum_{i}\hat{q}_{\nu}(i)$ $=\nu\sum_{i}\frac{p(i)}{g(i)}$
$=\nu c$
$<1,$
where we use that $\nu<\nicefrac{{1}}{{c}}$.
The next step is to prove that the information projection does not change if we restrict to the set of probability measures in $\mathcal{C}_{\nu^{*}}$, which we denote by $\tilde{\mathcal{C}}_{\nu^{*}}$. To this end, note first that for $\nu<\nu^{*}$, we have that $\sum g\left(i\right)q_{\nu}({i})<\nu^{*}$, so that for all $\nu<\nu^{*}$ there exists $n_{\nu}\in\mathbb{N}$ such that the probability measure defined by
$q_{\nu}(i)+\left(1-\sum_{j}q_{\nu}(j)\right)\delta_{n_{\nu}}\left(i\right)$
is an element of $\tilde{\mathcal{C}}_{\nu^{*}}$. Hence
$D(P\|\tilde{C})\leq D\left(P\left\|Q_{\lambda}+\left(1-\sum_{j}q_{\nu}({i})\right)\delta_{n_{\nu}}\right.\right)\\ =\sum_{i\in\mathbb{N}}p(i)\ln\left(\frac{p(i)}{Q_{\nu}({i})+\left(1-\sum_{j\in\mathbb{N}}q_{\nu}(j)\right)\delta_{n_{\nu}}\left(i\right)}\right)\\ =-p({n_{\nu}})\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})}\right)\\ +p(n_{\nu})\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})+1-\sum_{j\in\mathbb{N}}q_{\nu}({j})}\right)\\ +\sum_{i=1}^{\infty}p({i})\ln\left(\frac{p(i)}{q_{\nu}({i})}\right).$
The first term can be written as
$p({n_{\nu}})\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})}\right)$ $=q_{\nu}({n_{\nu}})\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})}\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})}\right)$
$=q_{\nu}({n_{\nu}})\frac{g\left(n_{\nu}\right)}{\nu}\ln\left(\frac{g\left(n_{\nu}\right)}{\nu}\right)$
Then notice that for $\nu\to\nu^{*}$, we must have that $n_{\nu}\to\infty$. Using that $c\ln\left(c\right)\to 0$ for $c\to 0$ we see the first term tends to 0 for $\nu\to\nu^{*}$.
Similarly, the second term can be written as
$p(n_{\nu})\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})+1-\sum_{j\in\mathbb{N}}q_{\nu}({j})}\right)=\\ \left(q_{\nu}(n_{\nu})+1-\sum_{j\in\mathbb{N}}q_{\nu}(j)\right)\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})+1-\sum_{j\in\mathbb{N}}q_{\nu}({j})}\\ \times\ln\left(\frac{p(n_{\nu})}{q_{\nu}({n_{\nu}})+1-\sum_{j\in\mathbb{N}}q_{\nu}({j})}\right).$
We also have
$\frac{p(n_{\nu})}{q_{\nu}(n_{\nu})+1-\sum q_{\nu}(i)}\to 0$
for $\nu\to\nu^{*}$ and using that $c\ln\left(c\right)\to 0$ for $c\to 0$ we get the second term tends to 0 for $\nu\to\nu^{*}$. Therefore we see
$D(P\|\tilde{\mathcal{C}})$ $\leq\lim_{\nu\to\nu^{*}}D\left(P\left\|Q_{\nu}+\left(1-\sum_{i}q_{\nu}({i})\right)\delta_{n_{\nu}}\right.\right)$
$\leq\sum_{i}p(i)\ln\left(\frac{p(i)}{q_{\nu^{*}}({i})}\right)$
$=\inf_{Q\in\mathcal{C}}\sum_{i}p(i)\ln\left(\frac{p(i)}{q(i)}\right).$
The inequality trivially also holds the other way around, so we find that
$D(P\|\tilde{\mathcal{C}})=\inf_{Q\in\mathcal{C}}\sum_{i}p(i)\ln\left(\frac{p(i)}{q(i)}\right).$
It follows that $Q_{\nu^{*}}$ is a strict sub-probability measure, and at the same time it is the reverse information projection of $P$ onto $\tilde{\mathcal{C}}_{\nu^{*}}$.